OpenStudy (anonymous):

which ordered pair is a solution of the equation 4x -2y=16

OpenStudy (jiteshmeghwal9):

2(2x-y)=16 2x-y=16/2 2x-y=8

OpenStudy (anonymous):

that dosent match any of the answers

OpenStudy (anonymous):

theres an infinite number of answers but all you need to do is plug the points in and see if the left hand side is equal to the right hand side to see if it is a particular solution

OpenStudy (anonymous):

thank you